OpenStudy (dragon123mama):

<1 and <2 are supplementary angels. m<1 is 4Y+7, and m<2 is 9Y+4. find M<2?

Vocaloid (vocaloid):

supplementary angles mean that <1 and <2 add up to 180 so 4Y + 7 + 9Y + 4 = 180 first step, solve for Y

OpenStudy (dragon123mama):

i cannot figure this problem out at all and i have tried to but i just cant.

OpenStudy (dragon123mama):

oh okay thanks, ill try that then.

OpenStudy (dragon123mama):

so 13Y+11=180

Vocaloid (vocaloid):

good so far, now find Y

OpenStudy (dragon123mama):

y=13

Vocaloid (vocaloid):

good. now <2 = 9Y + 4 = ?

OpenStudy (dragon123mama):

121

OpenStudy (dragon123mama):

am i right?

Vocaloid (vocaloid):

yup! that's our final answer

OpenStudy (dragon123mama):

ohhh okay this makes more sence now thank you